How many grams of zinc chloride could be formed from the reaction of 3.57 g of zinc with excess HCl?
1
Expert's answer
2020-11-11T10:36:46-0500

Zn+2HCl = ZnCl2 + H2

Zn= ZnCl2

65.38g Zn= 136.315 g ZnCl2

1g Zn = 136.315/65.38 g ZnCl2

3.57 g Zn = (136.315/ 65.38) × 3.57 g of ZnCl2

3.57g Zn = 7.443g ZnCl2